A 30.0 mL sample of 0.200 mol/L hydrobromic acid solution (HBr(aq)) was titrated with a 0.200 mol/L potassium hydroxide solution (KOH(aq)). Which one of the following statements concerning the pH curve of this reaction would be incorrect?
A. There would be one sharp change in pH during the titration.
B. The pH of the solution would increase as the titration progressed.
C. Methyl orange could be used to detect the point at which the reaction was complete.
D. The pH of the hydrobromic acid solution would change slowly at the beginning of the titration.
1
Expert's answer
2019-07-22T07:22:41-0400

Solution.

"HBr + KOH = KBr + H2O"

"H^+ + OH^- = H2O"

HBr is a strong acid.

KOH is a strong acid.

During the titration, we will observe one sharp change in pH. Since we are going to use a strong base as the titrant, when adding alkali, the pH of the solution will gradually increase. Since we titrate HBr, which is a strong acid, up to the equivalence point, the pH will change very slowly.


"C(HBr) \times V(HBr) = C(KOH) \times V(KOH)"

"V(KOH) = \frac{C(HBr) * V(HBr)}{C(KOH)}"

V(KOH) = 0.03 L = 30 ml

Alkali volume at 99.9% titration:


"V(KOH) = \frac{30 \times 99.9}{100}"

V(KOH) = 29.97 ml

The concentration of the not titrated amount of acid:


"C(HBr) = [H^+] = \frac{0.2 \times 0.03 - 0.2 \times 0.02997}{0.03+0.02997}"

C(HBr) = 0,00010005002 mole/L


"pH = -lg([H^+])"

pH = 4.00

At the point of the end of the titration step, the solution is overtight by 0.1%:


"V(KOH) = \frac{30 \times 100.1}{100}"

V(KOH) = 30.03 ml


"C(KOH) = [OH^-] = \frac{0.2 \times 0.03003 - 0.2 \times 0.03}{0.03+0.03003}"

C(KOH) = 0,00009995002 mole/L


"pOH = -lg([OH^-])"

pOH = 4.0002


"pH = 14 - pOH"

pH = 10.00

Thus, the titration jump will be in the region from 4.00 to 10.00. Methyl orange as an indicator is not suitable here, since the interval of color change for this indicator is from 3.1 to 4.4. In this case, you can use Bromthymol Blue or Phenolphthalein as an indicator.

Answer:

C